OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (4, 8)

OpenStudy (anonymous):

You can use the same method for the question you asked earlier, it was the same question with different x,y values. Use the same method and formula.

OpenStudy (anonymous):

First find slope(m) for the line . this line should come in this points : A=(7,2) B=(4,8) m=Y2-Y1/X2-X1 y2=2 y1=8 x2=4 x1=7 Then use this form : y2-y=m(x2=x) (x and y are not x1 and y1 in the first . they are in the form : y=mx+d ) @Annа :Did you get it ?;)

OpenStudy (the_fizicx99):

Its y-y1=m(x-x1)

OpenStudy (anonymous):

@tHe_FiZiCx99 :‌ Hi ! Are you sure ?!!!

OpenStudy (the_fizicx99):

Pretty sure